OpenStudy (anonymous):

Find the sum of 2 and the square root of 6 and the square root of 24

OpenStudy (anonymous):

1452

OpenStudy (anonymous):

ya ok

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i mean ure just gonna take anything any1 gives u no?

OpenStudy (anonymous):

no thinkin rite?

OpenStudy (anonymous):

lol hehe u have a good sense of humour pre algbra

OpenStudy (anonymous):

First off, learn to spell. Secondly, hop off.

OpenStudy (anonymous):

im russian u speak russian?

OpenStudy (anonymous):

privyet

OpenStudy (anonymous):

No, I'm not Russian. But, just because you are Russian, doesn't mean that you can't learn correct grammar.

OpenStudy (anonymous):

kagdyela?

OpenStudy (anonymous):

First - see if you can simplify the \[\sqrt{24}\] are there any square factors of 24?

OpenStudy (anonymous):

2 and 4.

OpenStudy (anonymous):

yeii nuzhno nauchit'sya uchit'sya

OpenStudy (anonymous):

so yes 4 is a square factor so you can write it also as \[\sqrt{4}*\sqrt{6}\]

OpenStudy (anonymous):

hehe dont know russian that well. My good friend was russian. so she taught me a lil. actually she is from ukraine she wld kill me if i called her russian

OpenStudy (anonymous):

what does that simplify to?

OpenStudy (anonymous):

2 and the square root of 30 ? or just stays 4 and the square root of 6 ?

OpenStudy (anonymous):

so the equation right now is \[2+\sqrt{6} + \sqrt{4}*\sqrt{6}\]

OpenStudy (anonymous):

what does \[\sqrt{4}\] equal?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

k - so we now have \[2+\sqrt{6}+2*\sqrt{6}\] can that be further simplified?

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

It can!

OpenStudy (anonymous):

4 and the square root of 6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nope. What does x+2x equal?

OpenStudy (anonymous):

2 and the square root of 30

OpenStudy (anonymous):

you've got similar terms, both with the sqrt(6) you need to combine those terms

OpenStudy (anonymous):

i honestly don't know then. i have no idea.

OpenStudy (anonymous):

what does x+2x equal?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so its 4 and the square root of 6. aka \[4\sqrt{6}\]

OpenStudy (anonymous):

no. x+2x = 3x

OpenStudy (anonymous):

\[2\sqrt{30}\]

OpenStudy (anonymous):

my final answer. i'm sooo done with this question.

OpenStudy (anonymous):

Where are you getting 30?